A pair of dice are rolled and the sum on their upturned faces is recorded. What is the probability that the sum showing is 8, given that one die is showing a 5?
I thought it would be 1/6 because there are 6 sides of the dice and it would have to land on a 3 in order for the sum to be 8

Answer by solver91311(24713)



Exactly correct. Think about it this way: If you have already rolled a 5, there are exactly six possible outcomes for the sum of the two upward faces, 5 + 1 = 6, 5 + 2 = 7, 5 + 3 = 8, 5 + 4 = 9, 5 + 5 = 10, and 5 + 6 = 11. Only one of those combinations is 8, so only one result out of six possibilities represents success.

John